# Thermodynamic Bounds on the Asymmetry of Cross-Correlations with Dynamical Activity and Entropy Production

### 01 Dec 2023 v1

@@ Abstract
Introduction.—Entropy production plays a pivotal role in nonequilibrium thermodynamics and statistical mechanics, serving as a fundamental quantity of interest [1]. The quest to establish connections between entropy production and measurable parameters has been a central focus in this field. Any manifestation of nonequilibrium should be connected with entropy production or dissipation. For instantce, maintaining instantaneous equilibrium necessitates infinitely slow driving, implying that entropy production accompanies any finite-time process. This has led to the notion of thermodynamic geometry [2, 3, 4, 5] and speed limits in terms of entropy production [6, 7, 8, 9]. In equilibrium, currents vanish, so nonvanishing currents are also a signature of nonequilibrium, leading to the existence of bounds on entropy production in terms of nonvanishing currents normalized by its variance, known as thermodynamic uncertainty relations [10, 11, 12]. Additionally, equilibrium conditions give rise to the principle of microscopic reversibility, implying the symmetry of cross-correlations. Consequently, corresponding asymmetry also serves as a distinguishing feature of nonequilibrium steady states [13], and is presumably associated with entropy production [14].
@@
The numerator $\alpha_{ab}$, i.e., the slope of cross-correlation asymmetry (also called stationary fluctuation oscillation in Ref. [18]), vanishes in equilibrium, so a nonzero asymmetry implies nonequilibrium. The slopes of auto-correlations, $D_{aa}$ and $D_{bb}$, are a measure of diffusion [15]. Explicitly, the slopes of cross-correlation asymmetry and average auto-correlation can be expressed as
@@
where $\mathcal{N}^{*}$ is the number of nonzero local activities. The derivation of Eq. (8) for unicyclic networks and Eq. (9) for general cases is deferred to the end of the Letter. We present numerical results to support Eq. (8) in Fig. 1(d), with a total of $10^{6}$ data points and three distinct topologies as in Figs. 1(a)-(c). Each data point is generated as follows: we draw nonzero transition rates $W_{mn}$’s from the uniform distribution, calculate the diagonals of the transition rate matrix, and randomly sample the observables $a_{m}$ and $b_{m}(m=1,\ldots,\mathcal{N})$ from the interval $[-1,1]$. Subsequently, we calculate the bound and $|\chi_{ab}|$ using Eqs. (3)–(7). The scatter plot clearly validates the first bound for four-state networks, and networks with a different number of states show qualitatively similar results. The three topologies are explicitly considered because it is inherently impossible to obtain non-fully-connected networks [Figs. 1(b)-(c)] through random sampling, but they are qualitatively different from each other. This is because the bound is discontinuous at $W_{mn}=0$, arising from the presence of $\gamma$. For example, assume that $\mathcal{T}_{mn}$’s are finite for any $mn\neq 13$ and $31$, and all $p^{\text{ss}}_{m}$’s are finite as well. If $W_{13}=W_{31}=0$, then $\gamma=\max\mathcal{T}_{mn}/\min_{mn\neq 13,31}\mathcal{T}_{mn}$. However, if $W_{13}\ll 1$ and $W_{31}$ is finite, then $\gamma=\max\mathcal{T}_{mn}/\mathcal{T}_{13}$, which can be arbitrarily large.
It can be seen from the derivation that bound (8) saturates in unicyclic networks when the transition rates are uniform, the observables form a regular polygon, and the system is close to equilibrium. This saturation condition is similar to that of the thermodynamic uncertainty relation [10]. In contrast, bound (9) is saturated in unicyclic networks when the transition rates are uniform, the observables form a regular polygon, and the number of states $\mathcal{N}$ approaches infinity. Under these conditions, the two bounds are asymptotically equivalent. Otherwise, either bound can be tighter depending on the parameters.
@@
$\frac{\mathcal{A}\alpha_{ab}^{2}}{(D_{a}+D_{b})^{2}/4}$ $=\frac{16\sum_{i}\mathcal{A}_{i,i+1}(\sum_{i}\Omega_{i,i+1})^{2}\mathcal{J}^{2}}{(\sum_{i}\mathcal{A}_{i,i+1}L_{i,i+1}^{2})^{2}}$ (11)
@@
which is equivalent to Eq. (8). In the derivation of the first inequality, we used the inequality
$\frac{\sum x_{i}}{(\sum x_{i}y_{i})^{2}}\leq\left(\frac{\max x_{i}}{\min x_{i}}\right)^{2}\frac{\sum x_{i}^{-1}}{(\sum y_{i})^{2}},$ (12)
@@
Since $\left[4n_{c}^{\prime}\tan\left(\pi/n_{c}^{\prime}\right)\right]^{-1}$ is a monotonically increasing function of $n_{c}^{\prime}$ $(n_{c}^{\prime}>2)$ and $\left[n_{c}\tanh\left(\mathcal{F}_{c}/2n_{c}\right)\right]^{-1}$ a decreasing function of $n_{c}$ $(n_{c}>0)$, we have
@@
where we have used $n_{c}^{\prime}\leq n_{c}\leq\mathcal{N}$, $\mathcal{N}\tanh(\mathcal{F}_{c}/2\mathcal{N})/[4\mathcal{N}\tan(\pi/\mathcal{N})]\leq\mathcal{F}_{c}/8\pi$ and $\sigma_{c}=\mathcal{F}_{c}\mathcal{J}_{c}$. For the definitions of $n_{c}^{\prime}$, $n_{c}$ and $\mathcal{C}^{*}$, please refer to Ref. [15]. The equality holds for unicyclic systems when the observables form a regular polygon in the large $\mathcal{N}$ limit.
On the other hand, the normalized asymmetry is also bounded by
$\left|\chi_{ba}\right|$ $\leq\frac{4\sum_{c\in\mathcal{C}^{*}}\mathcal{J}_{c}\left|\sum_{e\in c}\Omega_{e}\right|}{\sum_{c\in\mathcal{C}^{*}}\sum_{e\in c}\mathcal{A}_{e}L_{e}^{2}}$ (16)
$\leq\frac{4\sum_{c\in\mathcal{C}^{*}}\mathcal{J}_{c}\left(\sum_{e\in c}L_{e}\right)^{2}\left[4n_{c}^{\prime}\tan\left(\pi/n_{c}^{\prime}\right)\right]^{-1}}{\min\mathcal{A}_{mn}\sum_{c\in\mathcal{C}^{*}}\sum_{e\in c}L_{e}^{2}}$
@@
where we have used the monotonicity of $\left[4n_{c}^{\prime}\tan\left(\pi/n_{c}^{\prime}\right)\right]^{-1}$, the isoperimetric inequality and Cauchy-Schwarz inequality again. The equality holds for unicyclic systems when the observables form a regular polygon.
Multiplying these two bounds on $|\chi_{ab}|$ and employing $\mathcal{A}\leq\mathcal{N}^{*}\max\mathcal{A}_{mn}$ results in
@@
Discussion.— In summary, we report two thermodynamic bounds [Eqs. (8) and (9)] on the normalized asymmetry of cross-correlation in terms of the dynamical activity and entropy production. Identifying a simpler expression that directly relates these three quantities seems to be a challenge. This bound exhibits broad applicability, regardless of the underlying network topology, and offers experimental testability. We also anticipate that this bound will provide valuable insights in addressing the conjecture proposed by Oberreiter et al. [17]. Rigorously establishing the validity of Eq. (8) for arbitrary network topologies does not appear to be straightforward, but to this end, the utilization of uniform cycle decomposition may serve as a crucial technique [44, 15]. A future research direction entails investigating whether this bound holds true or how it should be modified when applied to partially observed Markov networks [45, 46, 31]. Additionally, exploring the influence of quantum effects, such as quantum coherence, on this relation represents an intriguing avenue for further investigation.

### 05 Apr 2024 v2

@@ Abstract
Introduction.—Entropy production plays a pivotal role in nonequilibrium thermodynamics and statistical mechanics, serving as a fundamental quantity of interest [1]. The quest to establish connections between entropy production and measurable parameters has been a central focus in this field. Any manifestation of nonequilibrium should be connected with entropy production or dissipation. For instantce, maintaining instantaneous equilibrium necessitates infinitely slow driving, implying that entropy production accompanies any finite-time process. This has led to the notion of thermodynamic geometry [2, 3, 4, 5] and speed limits in terms of entropy production [6, 7, 8, 9, 10, 11]. In equilibrium, currents vanish, so nonvanishing currents are also a signature of nonequilibrium, leading to the existence of bounds on entropy production in terms of nonvanishing currents normalized by its variance, known as thermodynamic uncertainty relations [12, 13, 14]. Additionally, equilibrium conditions give rise to the principle of microscopic reversibility, implying the symmetry of cross-correlations. Consequently, corresponding asymmetry also serves as a distinguishing feature of nonequilibrium steady states [15], and is presumably associated with entropy production [16].
@@
The numerator $\alpha_{ab}$, i.e., the slope of cross-correlation asymmetry at $\tau=0$ (also called stationary fluctuation oscillation in Ref. [20]), vanishes in equilibrium, so a nonzero asymmetry implies nonequilibrium. The slopes of auto-correlations, $D_{aa}$ and $D_{bb}$, are a measure of diffusion [17]. Explicitly, the slopes of cross-correlation asymmetry and average auto-correlation can be expressed as
@@
where $\mathcal{N}^{*}$ is the number of nonzero local activities. The derivation of Eq. (8) for unicyclic networks and Eq. (9) for general cases is deferred to the end of the Letter. We conjecture that Eq. (8) is valid for arbitrary networks, and present numerical evidence in Fig. 1(d), with a total of $10^{6}$ data points and three distinct topologies as in Figs. 1(a)-(c). Each data point is generated as follows: We choose a topology randomly from the three topologies, draw nonzero transition rates $W_{mn}$’s from the uniform distribution, calculate the diagonals of the transition rate matrix, and randomly sample each of the observables $a_{m}$ and $b_{m}(m=1,\ldots,\mathcal{N})$ from the interval $[-1,1]$. Subsequently, we calculate the bound and $|\chi_{ab}|$ using Eqs. (3)–(7). The three topologies are explicitly considered because it is inherently impossible to obtain non-fully-connected networks [Figs. 1(b)-(c)] through random sampling, but they are qualitatively different from each other. This is because the bound is discontinuous at $W_{mn}=0$, arising from the presence of $\gamma$. For example, assume that $\mathcal{T}_{mn}$’s are finite for any $mn\neq 13$ and $31$, and all $p^{\text{ss}}_{m}$’s are finite as well. If $W_{13}=W_{31}=0$, then $\gamma=\max\mathcal{T}_{mn}/\min_{mn\neq 13,31}\mathcal{T}_{mn}$. However, if $W_{13}\ll 1$ and $W_{31}$ is finite, then $\gamma=\max\mathcal{T}_{mn}/\mathcal{T}_{13}$, which can be arbitrarily large. The scatter plot clearly validates the first bound for four-state networks, and networks with a different number of states show qualitatively similar results.
It can be seen from the derivation that bound (8) saturates in unicyclic networks when the transition rates are uniform, the observables form a regular polygon, and the system is close to equilibrium. This saturation condition is similar to that of the thermodynamic uncertainty relation [12]. In contrast, bound (9) is saturated in unicyclic networks when the transition rates are uniform, the observables form a regular polygon, and the number of states $\mathcal{N}$ approaches infinity. For unicyclic networks with uniform transition rates, we have $\mathcal{N}^{*}=\mathcal{N}$ and $\gamma=1$, and it can be proved that the first bound is tighter. As $\mathcal{N}\to\infty$, the two bounds are asymptotically equivalent. Otherwise, either bound can be tighter depending on the parameters.
@@
As proved in Ref. [17], the ratio between the real and imaginary components of the eigenvalue pertaining to the transition rate matrix can be regarded as a specific instance of the normalized asymmetry of cross-correlation. Thus, the two bounds also provide insight into the spectra of transition rate matrix from a thermodynamic standpoint, in line with ongoing research along this direction [18, 45, 19, 46, 47] . They could also contribute to the final resolution of the Oberreiter-Seifert-Barato conjecture [19] as the two bounds directly incorporate the entropy production (in comparison to [17]) and are observable-independent (in comparison to [20]).
@@
$\frac{\mathcal{A}\alpha_{ab}^{2}}{(D_{a}+D_{b})^{2}/4}$ $=\frac{16\sum_{i}\mathcal{A}_{i,i+1}(\sum_{j}\Omega_{j,j+1})^{2}\mathcal{J}^{2}}{(\sum_{i}\mathcal{A}_{i,i+1}L_{i,i+1}^{2})^{2}}$ (11)
@@
which is equivalent to Eq. (8). Note that for this case $\gamma=\max\mathcal{A}_{i,i+1}/\min\mathcal{A}_{i,i+1}$ as defined in Eq. (3). In the derivation of the first inequality, we used the inequality
$\frac{\sum x_{i}}{(\sum x_{i}y_{i})^{2}}\leq\frac{\sum[(\max\bm{x})^{2}/x_{i}]}{(\min\bm{x}\sum y_{i})^{2}}\leq\left(\frac{\max\bm{x}}{\min\bm{x}}\right)^{2}\frac{\sum x_{i}^{-1}}{(\sum y_{i})^{2}},$ (12)
@@
Here, $\mathcal{C}^{*}$ is the set of cycles with nonzero net asymmetry generated by the uniform cycle decomposition [49]. Given a cycle $c$, $n_{c}$ is the number of states, $n_{c}^{\prime}$ is the number of times the joint value $(a,b)$ changes throughout the duration of the cycle, $\mathcal{J}_{c}$ is the cycle current, and $\mathcal{F}_{c}$ is the cycle affinity. For more details of $n_{c}^{\prime}$, $n_{c}$ and $\mathcal{C}^{*}$, please refer to Ref. [17]. Since $\left[4n_{c}^{\prime}\tan\left(\pi/n_{c}^{\prime}\right)\right]^{-1}$ is a monotonically increasing function of $n_{c}^{\prime}$ $(n_{c}^{\prime}>2)$ and $\left[n_{c}\tanh\left(\mathcal{F}_{c}/2n_{c}\right)\right]^{-1}$ a decreasing function of $n_{c}$ $(n_{c}>0)$, we have
@@
where we have used $n_{c}^{\prime}\leq n_{c}\leq\mathcal{N}$, $\mathcal{N}\tanh(\mathcal{F}_{c}/2\mathcal{N})/[4\mathcal{N}\tan(\pi/\mathcal{N})]\leq\mathcal{F}_{c}/8\pi$ and $\sigma_{c}=\mathcal{F}_{c}\mathcal{J}_{c}$. The equality holds for unicyclic systems when the observables form a regular polygon in the large $\mathcal{N}$ limit.
On the other hand, one can use Eq. (S10) in Ref. [17] and the fact that $\mathcal{C}^{*}$ is a restricted set to bound the normalized asymmetry by
$\left|\chi_{ba}\right|\leq\frac{4\sum_{c\in\mathcal{C}^{*}}\mathcal{J}_{c}\left|\sum_{e\in c}\Omega_{e}\right|}{\sum_{c\in\mathcal{C}^{*}}\sum_{e\in c}\mathcal{A}_{e}L_{e}^{2}}.$ (16)
Furthermore, we obtain
$\left|\chi_{ba}\right|$ $\leq\frac{4\sum_{c\in\mathcal{C}^{*}}\mathcal{J}_{c}\left(\sum_{e\in c}L_{e}\right)^{2}\left[4n_{c}^{\prime}\tan\left(\pi/n_{c}^{\prime}\right)\right]^{-1}}{\min\mathcal{A}_{mn}\sum_{c\in\mathcal{C}^{*}}\sum_{e\in c}L_{e}^{2}}$ (17)
@@
where we have used the isoperimetric inequality, the monotonicity of $\left[4n_{c}^{\prime}\tan\left(\pi/n_{c}^{\prime}\right)\right]^{-1}$ and Cauchy-Schwarz inequality again. The equality holds for unicyclic systems when the observables form a regular polygon.
Multiplying these two bounds on $|\chi_{ab}|$ [Eqs. (15) and (17)] and employing $\mathcal{A}\leq\mathcal{N}^{*}\max\mathcal{A}_{mn}$ results in
@@
Discussion.— In summary, we report two thermodynamic bounds [Eqs. (8) and (9)] on the normalized asymmetry of cross-correlation in terms of the dynamical activity and entropy production. Identifying a simpler expression that directly relates these three quantities seems to be a challenge. This bound exhibits broad applicability, regardless of the underlying network topology, and offers experimental testability. We also anticipate that this bound will provide valuable insights in addressing the conjecture proposed by Oberreiter et al. [19]. Rigorously establishing the validity of Eq. (8) for arbitrary network topologies does not appear to be straightforward, but to this end, the utilization of uniform cycle decomposition may serve as a crucial technique [49, 17]. In view of the Langevin equation being regarded as the continuous-space limit of the master equation, we anticipate that our two bounds hold for overdamped Langevin systems, while additional terms might be needed for underdamped cases [50, 51]. A future research direction entails investigating whether this bound holds true or how it should be modified when applied to partially observed Markov networks [52, 53, 33]. Additionally, exploring the influence of quantum effects, such as quantum coherence, on this relation represents an intriguing avenue for further investigation.
Acknowledgment.— We thank Kangqiao Liu for valuable discussions.